OpenStudy (anonymous):

i can't figure out the answer 2p-3t=4 3p+2t=6 (solving systems by using elimination)

OpenStudy (anonymous):

Eliminate one of the variables (2p-3t=4) x 2 (3p+2t=6) x 3 4p-6t=8 9p+6t=18 Get rid of the t 13p=26 p=2 Plug back into the equation 2(2)-3t=4 4-3t=4 -3t=0 t=0 Check on other equation 3(2)+2(t)=6 6+2t=6 2t=0 t=0 P=2, T=0